Sommer, Inc., is considering a project that will result in initial aftertax cash savings of $1.78 million at the end of the first year, and these savings will grow at a rate of 2 percent per year indefinitely. The firm has a target debt-equity ratio of .80, a cost of equity of 11.8 percent, and an aftertax cost of debt of 4.6 percent. The cost-saving proposal is somewhat riskier than the usual project the firm undertakes; management uses the subjective approach and applies an adjustment factor of 3 percent to the cost of capital for such risky projects.

What is the maximum initial cost the company would be willing to pay for the project? (Do not round intermediate calculations and enter your answer in dollars, not millions, rounded to the nearest whole dollar amount, e.g., 1,234,567.)
Answer #1

The maximum initial cost the company would be willing to pay for the project is computed as shown below:

D/E = 0.80

Debt / Capital is computed as shown below:

= 0.80 / 1.80

= 0.4444

Equity / Capital is computed as follows:

= 1 / 1.80

= 0.5556

Cost of debt = 4.6% or 0.046

Cost of equity = 11.8% or 0.118

WACC will be as follows:

= 0.4444 x 0.046 + 0.5556 x 0.118

= 8.60%

Project discount rate will be as follows:

= WACC + Adjustment factor

= 8.60% + 3%

= 11.60%

PV will be as follows:

= $ 1.78 million / ( 0.1160 - 0.02 )

= $ 18,541,666.67

Feel free to ask in case of any query relating to this question